# EFKAN: A KAN-Integrated Neural Operator For Efficient Magnetotelluric Forward Modeling

### v1

@@ Abstract
Magnetotelluric (MT) forward modeling is fundamental for improving the accuracy and efficiency of MT inversion. Neural operators (NOs) have been effectively used for rapid MT forward modeling, demonstrating their promising performance in solving the MT forward modeling-related partial differential equations (PDEs). Particularly, they can obtain the electromagnetic field at arbitrary locations and frequencies. In these NOs, the projection layers have been dominated by multi-layer perceptrons (MLPs), which may potentially reduce the accuracy of solution due to they usually suffer from the disadvantages of MLPs, such as lack of interpretability, overfitting, and so on. Therefore, to improve the accuracy of MT forward modeling with NOs and explore the potential alternatives to MLPs, we propose a novel neural operator by extending the Fourier neural operator (FNO) with Kolmogorov-Arnold network (EFKAN). Within the EFKAN framework, the FNO serves as the branch network to calculate the apparent resistivity and phase from the resistivity model in the frequency domain. Meanwhile, the KAN acts as the trunk network to project the resistivity and phase, determined by the FNO, to the desired locations and frequencies. Experimental results demonstrate that the proposed method not only achieves higher accuracy in obtaining apparent resistivity and phase compared to the NO equipped with MLPs at the desired frequencies and locations but also outperforms traditional numerical methods in terms of computational speed.
@@ 1 Introduction
Magnetotelluric (MT) has been widely used to illuminate the deep earth by inverting the subsurface resistivity. This method utilizes the natural electromagnetic fields induced by solar and ionospheric currents, which interact with the resistivity of the earth to produce secondary fields measurable at the surface. Since its inception, MT has played a crucial role in various applications, including natural resource exploration (Jiang et al., 2022), geothermal energy assessment (Cheng et al., 2022), environmental monitoring (Romano et al., 2014), and deep geological structure characterization (Egbert et al., 2022).
The interpretation of MT data is grounded in the inverse problem of determining the resistivity of the earth from the observed electromagnetic field variations. The effectiveness and efficiency of MT inversion are usually significantly determined by forward modeling. Over the last decades, numerical computational methods such as finite difference method (FDM) (Varilsuha and Candansayar, 2018), finite element method (FEM) (Zhu et al., 2022), and finite volume method (FVM) (Guo et al., 2020) have dominated the solution of MT forward-related PDEs. These methods solve the PDEs with physical laws, thus they can obtain reasonable solutions. However, their accuracy largely depends on the spatial discretization; finer mesh grids yield higher accuracy but also increase computational costs.
Deep neural networks (DNNs) have been used for rapid MT forward modeling. This approach aims to learn the mapping function between the resistivity and the responses on the surface, such as the electromagnetic field or apparent resistivity, by training DNNs with supervision. Conway et al. (2019) trained a DNN that can output apparent resistivity and phase for three-dimensional (3-D) MT forward modeling. Shan et al. (2022) applied multitask learning (that is, a DNN has two branches) to predict the apparent resistivity and the phase of the mode $xy$. Wang et al. (2024) employed the 3-D Swin Transformer (Liu et al., 2021) for 3-D MT forward modeling, predicting the apparent resistivity and phase in different polarization directions. One main disadvantage of these methods is that their efficiency is limited by the spatial discretization of the resistivity model, meaning that the trained DNN only works for specific resolutions and frequencies. Thus, it is difficult to achieve satisfactory results on meshes or at frequencies that differ from the training data.
@@
However, FNOs have limitations in MT forward modeling because they require the output resolution to match the input resolution, which restricts their ability to obtain responses at arbitrary locations and frequencies. To address this shortcoming, DeepONets introduced the trunk net to map the output of the branch net to the desired solution based on the universal approximation theorem of operators (Lu et al., 2021). Peng et al. (2022) proposed the extended Fourier neural operator (EFNO) for MT forward modeling in the frequency domain by extending FNO with MLP. The architecture of EFNO is well suited for the demands of responses at arbitrary locations and frequencies for MT forward modeling.
As an alternative to MLPs, KANs, as proposed by Liu et al. (2024), address the shortcomings of MLPs, including a lack of interpretability (Cranmer, 2023) and overfitting. Unlike traditional NNs that use fixed activation functions at the nodes, KANs introduce learnable activation functions on the edges, which connect the nodes. These activation functions are typically parameterized by certain functions, such as spline functions, providing a high degree of flexibility and adaptability to model complex relationships with potentially fewer parameters. KANs have been explored for solving PDEs and operator learning. Liu et al. (2024) combined PINNs with KANs to solve the 2-D Poisson equation. Abueidda et al. (2024) proposed a radial basis functions (RBF)-based KAN operator for solving orthotropic elasticity problems. Shukla et al. (2024) conducted a comprehensive and fair comparison between MLP and KAN for PDE and operator learning.
In this study, we exploit the potential of KANs to improve the accuracy of MT forward modeling by developing a novel approach that leverages the advantages of FNOs, DeepONets, and KANs. Specifically, based on the basic framework of DeepONets, we utilize FNO as the branch network to handle the frequency information of the resistivity model, and we take the KAN as the trunk net to obtain the apparent resistivity and phase at the desired locations and frequencies. Furthermore, we employ the spectral method to generate the resistivity model with anomalies, rather than simply embedding anomalies within homogeneous half-space underground to evaluate the effectiveness of the proposed method. This paper is organized as follows. In Section 2, we set up the problem formulation, briefly describe neural operators and KANs, and then provide details about the proposed method. In Section 3, we demonstrate the effectiveness of the proposed method through numerical experiments. In Section 4, we provide a discussion about the characteristics of the proposed method. Finally, we draw our conclusions in Section 5.
@@ 2 Methodology
In this section, we first set up the MT forward problem with a surrogate model, followed by describing the proposed approach comprehensively.
@@ 2.1 Problem Formulation
In the frequency domain, the source-free propagation of the natural electric field $E$ and magnetic field $H$ can be expressed by the simplified Maxwell’s equations,
{×$\left\{\begin{array}[]{l}\nabla\times H=(\sigma-i\omega\varepsilon)E,\\ \nabla\times E=i\omega\mu H,\end{array}\right.$ (1)
where $\sigma(S/m)$ represents the conductivity of the subsurface medium, which is the reciprocal of resistivity $\rho(\Omega\cdot m)$, $i$ is the imaginary unit, $\varepsilon(F/m)$ denotes the permittivity, $\omega=2\pi f$ indicates the angular frequency of the electromagnetic (EM) field, and $\mu(H/m)$ denotes the permeability. For MT, it is commonly assumed that the permittivity and permeability of the earth are constants, e.i., $\varepsilon_{0}=8.85\times 10^{-12}\mathrm{~{}F}/\mathrm{m}$ and $\mu_{0}=4\pi\times 10^{-7}\mathrm{H}/\mathrm{m}$. Due to $\omega\varepsilon\ll\sigma$, the displacement current can be neglected. Therefore, the EM fields studied by the EM equations will satisfy the quasi-stationary field assumption within the frequency range of MT, and Equation 1 can be rewritten as
$\begin{cases}\nabla^{2}E-k^{2}E=0,\\ \nabla^{2}H-k^{2}H=0,\end{cases}$ (2)
where $k=\sqrt{-iw\mu\sigma}$. Thus, the number of PDEs to be solved can be reduced to those related to either $E$ or $H$, instead of solving all PDEs concerning both $E$ and $H$. For 2-D MT forward modeling, the conductivity varies only in the vertical direction $z$ and the horizontal direction $y$; therefore, we can decouple Equation 2 into two modes: mode $xy(E_{x},H_{y},H_{z})$ and mode $yx(H_{x},E_{y},E_{z})$ by
@@
where $u_{bc}$ presents the electric or magnetic field at the boundary, and $g(y,z)$ denotes the Dirichlet boundary condition at coordinates $y$ and $z$. Once we obtain $E$ and $H$, we can determine the apparent resistivity $\rho_{xy}$ and $\rho_{yx}$, as well as the phase $\phi_{xy}$ and $\phi_{yx}$, using
@@
where Re denotes the real part and Im denotes the imaginary part, respectively. Additionally, our primary concern is the electromagnetic field on the earth surface, as this is where we are typically able to deploy instruments for the collection of electromagnetic field data. To obtain the apparent resistivity and phase on the surface, conventional numerical modeling methods, such as the FDM, are often used to solve Equation 3, which can be represented as
@@
where $G_{\theta}$ indicates the solver with parameter $\theta$ for the desired apparent resistivity and phase. Classical linear MT forward modeling method, such as FDM, usually require discretizing the conductivity, followed by constructing linear systems for different frequencies. Hence, the computational cost of conventional methods is very high, as their performance depends on the number of grids, and they require constructing a separate operator for each frequency. Therefore, it is necessary to develop a surrogate model of $G_{\theta}$ to improve the efficiency of MT forward modeling.
@@ 2.2.1 Neural Operators
Neural operators have been used to solve PDEs with modern machine learning technique, which the solution operator. Representative models are FNOs (Li et al., 2020) and DeepONets (Lu et al., 2021).
To construct a NN for solving PDEs, Li et al. (2020) proposed the NO, which can be written as
@@
where $v_{j+1}$ denotes the output of the $j+1$-th layer of the NN, $\alpha$ represents the activation function, $v_{j}$ indicates the output from the previous layer, and $W$ is the weight. Compared to the standard NN, the right-hand side of Equation 7 includes an additional integral term, where $\kappa$ signifies the integral kernel function that needs to be learned from data. By incorporating this integral term, the NO can extract non-local features from the data. Due to the fact that the computation of the integral is time-consuming, Li et al. (2021) further proposed the FNO by Fourier layer that transforms the integral operation of Equation 7 into frequency domain. The Fourier layer can be expressed as
@@
where $\mathcal{F}$ denotes the Fourier transform, and $\mathcal{F}^{-1}$ is the inverse Fourier transform. As shown in Fig. 1, FNO consists of lifting layer $P$, a Fourier layer, and projecting layer $Q$. The lifting layer $P$ maps the input $a$ into a high-dimensional channel space. The Fourier layer transforms the input $v$ into frequency and filtering the Fourier coefficients, followed by inverse Fourier transform. The $W$ represents linear transformations, such as those performed by MLPs. The projecting layer $Q$ projects the data to the desired dimension.
Although FNO can share the same parameters regardless of the discretization employed for both input and output (Kovachki et al., 2024), it is not suitable for MT forward modeling. This is because FNO typically requires that the input and output data have consistent spatial dimensions. This means that if the input data is a multi-dimensional array with a specific resolution, then the output data should also have the same dimensions and resolution. For MT forward modeling, the results of the forward modeling depend not only on the resistivity model but also on the observation frequency and the location of the electromagnetic field observation.
Unlike FNO, DeepONet is not restricted to a specific architecture and can incorporate various types of neural network architectures in its branch and trunk nets, providing greater flexibility. The formulation of DeepONet can be expressed as
@@
where $b_{k}$ represents $k$-th branch net, $t_{k}$ denotes $k$-th trunk net, and $p$ is the number of branches. In DeepONet, branch and trunk nets facilitate the approximation of infinite-dimensional mappings from inputs to outputs, offering a highly flexible and efficient approach. The branch network provides the necessary input features, while the trunk network maps these features to the output space, enabling DeepONet to handle complex and high-dimensional operator learning tasks (Tianping Chen and Hong Chen, 1995). This architecture allows DeepONet to generalize well across a range of scenarios, including different domain geometries, input parameters, and initial and boundary conditions (Kontolati et al., 2024).
@@ 2.2.2 KANs
In KANs, the activation functions are not fixed; instead, they are learned during the training process, allowing for a more flexible and efficient approximation of complex functions. Each weight parameter in a KAN is replaced with a univariate function. In the original implementation (Liu et al., 2024), $\psi(x)$ is defined as a weighted combination of a basis function $b(x)$ and B-splines,
@@
where $b(x)$ and spline$(x)$ are defined as follows,
@@
where $w_{b}$,$w_{s}$ and $c_{i}$ are the weights that are optimized during training, and $B_{i}(x)$ are the B-spline basis functions defined over a grid.
Thus, the proposed method is able to improve the interpretability, as the learnable functions can be visualized and understood more intuitively. The training process involves optimizing the weights of these splines to minimize the loss function, adjusting the shape of the spline to best fit the training data. KANs offer a promising alternative to traditional MLPs by leveraging the theoretical foundation of the Kolmogorov-Arnold representation theorem, leading to potentially more efficient and interpretable models for approximating complex functions.
@@ 2.2.3 EFKAN
Considering the characteristics of MT forward modeling, as well as the advantages of FNOs, DeepONets, and KANs, we propose extending FNO with KAN to improve the accuracy of MT forward modeling, which we name EFKAN (Fig. 2). Specifically, we employ FNO as the branch net to map the resistivity $\rho(y,z)$ to $U(y,z)$; for the trunk net, we utilize the desired frequencies and coordinates of the electromagnetic field observation as inputs to KAN to obtain $D(y,z,f)$. Note that $D(y,z,f)$ has the same dimension with $U(y,z)$. We can therefore achieve the desired apparent resistivity and phase ($\rho_{xy},\phi_{xy},\rho_{yx},\phi_{yx}$) by the matrix multiplication of $u_{i}$ and $d_{i}$. For the 2-D case, the mathematical expression of MT forward modeling with EFKAN can be written as
@@
In the FNO of EFKAN, we utilize the fast Fourier transformer (FFT), 2-D convolution, and GELU to construct the Fourier layer. In the frequency domain, it is necessary to truncate the high modes to obtain $\mathcal{F}\left(v_{t}\right)\in\mathbb{C}^{k_{\max}\times d_{v}}$ because we convolve $v_{t}\in\mathbb{R}^{n\times d_{v}}$ with a function that has only $k_{\max}$ Fourier modes. We set $k_{\max}$ to 18. For the 2-D convolution, we define the kernel size to $1\times 1$, and keep the input channels equal to the output channels. Additionally, we place linear layers at the beginning and end of the FNO to determine the dimension of output, and we employ the Gaussian error linear unit (GELU) as the activation function. More details related to the FNO can be found in Table 1. For the KAN of EFKAN, we use the original KAN implemented in (Liu et al., 2024) that adopts B-spline function as the activation function. The trunk net-KAN has three layers, and the details about each layer are listed in Table 2.

### v2

@@ Abstract
Forward modeling is the cornerstone of magnetotelluric (MT) inversion. Neural operators have been successfully applied to solve partial differential equations, demonstrating encouraging performance in rapid MT forward modeling. In particular, they can obtain the electromagnetic field at arbitrary locations and frequencies, which is meaningful for MT forward modeling. In conventional neural operators, the projection layers have been dominated by classical multi-layer perceptrons, which may reduce the precision of solution because they usually suffer from the disadvantages of multi-layer perceptrons, such as lack of interpretability, overfitting, etc. Therefore, to improve the accuracy of the MT forward modeling with neural operators, we integrate the Fourier neural operator with the Kolmogorov-Arnold network (KAN). Specifically, we adopt KAN as the trunk network instead of the classic multi-layer perceptrons to project the resistivity and phase, determined by the branch network-Fourier neural operator, to the desired locations and frequencies. Experimental results demonstrate that the proposed method can achieve high precision in obtaining apparent resistivity and phase at arbitrary frequencies and/or locations with rapid computational speed.
@@ 1 Introduction
Magnetotelluric (MT) has been widely used to illuminate the deep Earth by inverting the sampled electromagnetic field on the Earth surface. MT utilizes natural electromagnetic field induced by solar and ionospheric currents, which interact with the resistivity of the earth to produce secondary fields measurable at the surface. MT has played a crucial role in various applications, including exploration of natural resources (Jiang et al., 2022; Smith, 2014; Wang et al., 2024b), geothermal energy assessment (Cheng et al., 2022; Gao et al., 2018), environmental monitoring (Romano et al., 2014), and characterization of deep geological structures (Egbert et al., 2022; Yu et al., 2020, 2024a).
The interpretation of MT data is grounded in the inverse problem of determining the resistivity of the Earth from the observed electromagnetic field variations. The classic iterative MT inversion methods (e.g. nonlinear conjugate gradient (Newman and Alumbaugh, 2000) and OCCAM (Siripunvaraporn and Egbert, 2009)) consist of forward modeling and model optimization. The essence of forward modeling is to solve the Helmholtz equation for a given conductivity model. The model optimization is to search for the optimal solution by minimizing the mismatch between the observation and the forward modeling electromagnetic field. The efficiency of the MT inversion is usually significantly determined by forward modeling, especially for the high-dimensional MT data inversion. In recent decades, numerical computational methods such as the finite difference method (FDM) (Varilsuha and Candansayar, 2018), finite element method (Zhu et al., 2022), and the finite volume method (Guo et al., 2020) have dominated the solution of forward related PDEs of MT. These methods solve the PDEs with physical laws and thus can obtain reasonable solutions. However, their precision largely depends on spatial discretization; finer mesh grids yield higher accuracy but also increase computational costs.
Deep neural networks (DNNs) are powerful for rapid MT forward modeling, which can directly map the conductivity to the observation (e.g. electromagnetic field or apparent resistivity) on the Earth surface. Conway et al. (2019) trained a DNN that can output apparent resistivity and phase for three-dimensional (3-D) MT forward modeling. Shan et al. (2022) applied multitask learning (that is, a DNN has two branches) to predict the apparent resistivity and the phase of the mode $xy$. Wang et al. (2024a) employed the 3-D Swin Transformer (Liu et al., 2021) for 3-D MT forward modeling, predicting the apparent resistivity and phase in different polarization directions. One main disadvantage of these methods is that their efficiency is limited by spatial discretization of the resistivity model, meaning that the trained DNN only works for specific resolutions and frequencies. Thus, it is difficult to achieve satisfactory results on meshes or at frequencies that differ from the training data.
@@
However, FNOs have limitations in MT forward modeling because they require the output resolution to match the input resolution, restricting their ability to obtain responses at arbitrary locations and frequencies. To address this shortcoming, DeepONets introduced the trunk net to map the output of the branch net to the desired solution based on the universal approximation theorem of operators (Lu et al., 2021). Peng et al. (2022) proposed the extended Fourier neural operator (EFNO) for MT forward modeling in the frequency domain by extending FNO with multilayer perceptron (MLP). The architecture of EFNO is well suited for the demands of responses at arbitrary locations and frequencies for MT forward modeling.
As an alternative to multi-layer perceptrons (MLPs), KANs, as proposed by Liu et al. (2024), address the shortcomings of MLPs, including lack of interpretability (Cranmer, 2023) and overfitting. Unlike traditional NNs that use fixed activation functions at the nodes, KANs introduce learnable activation functions at the edges that connect the nodes. These activation functions are typically parameterized by certain functions, such as spline functions, providing a high degree of flexibility and adaptability to model complex relationships with potentially fewer parameters. KANs have been explored for solving PDEs and operator learning. Liu et al. (2024) combined PINNs with KANs to solve the 2-D Poisson equation. Abueidda et al. (2024) proposed a radial basis functions (RBF)-based KAN operator to solve orthotropic elasticity problems. Shukla et al. (2024) conducted a comprehensive and fair comparison between MLP and KAN for PDE and operator learning.
In this study, we exploit the potential of KANs to improve the accuracy of MT forward modeling by developing a novel approach that takes advantage of FNOs, DeepONets, and KANs. Specifically, based on the basic framework of DeepONets, we utilize FNO as the branch network to handle the frequency information of the resistivity model, and we take the KAN as the trunk net to obtain the apparent resistivity and phase at the desired locations and frequencies. Furthermore, we employ the spectral method to generate the resistivity model with anomalies, rather than simply embedding anomalies within homogeneous half-space underground to evaluate the effectiveness of the proposed method. This paper is organized as follows. In section 2, we set up the problem formulation, briefly describe neural operators and KANs, and then provide details about the proposed method. In section 3, we demonstrate the effectiveness of the proposed method through numerical experiments. In section 4, we provide a discussion of the characteristics of the proposed method. Finally, we draw our conclusions in section 5.
@@ 2 Methodology
In this section, we first establish the MT forward problem with a surrogate model, followed by describing the proposed approach.
@@ 2.1 Problem Formulation
In the frequency domain, source-free propagation of the natural electric field E and the magnetic field H can be expressed using the simplified Maxwell equations,
{×$\left\{\begin{array}[]{l}\nabla\times\textbf{H}=(\sigma-i\omega\varepsilon)\textbf{E},\\ \nabla\times\textbf{E}=i\omega\mu\textbf{H},\end{array}\right.$ (1)
where $\sigma(S/m)$ represents the conductivity of the subsurface medium, which is the reciprocal of resistivity $\rho(\Omega\cdot m)$, $i$ is the imaginary unit, $\varepsilon(F/m)$ denotes the permittivity, $\omega=2\pi f$ indicates the angular frequency of the electromagnetic field, and $\mu(H/m)$ denotes the permeability. For MT, it is commonly assumed that the permittivity and permeability of the earth are constants, e.i. $\varepsilon_{0}=8.85\times 10^{-12}\mathrm{~{}F}/\mathrm{m}$ and $\mu_{0}=4\pi\times 10^{-7}\mathrm{H}/\mathrm{m}$. Due to $\omega\varepsilon\ll\sigma$, the displacement current can be neglected. Therefore, the electromagnetic field studied by the electromagnetic equations will satisfy the quasi-stationary field assumption within the frequency range of MT, and the equation 1 can be rewritten as
$\begin{cases}\nabla^{2}\textbf{E}-k^{2}\textbf{E}=0,\\ \nabla^{2}\textbf{H}-k^{2}\textbf{H}=0,\end{cases}$ (2)
where $k=\sqrt{-iw\mu\sigma}$. Thus, the number of PDEs to be solved can be reduced to those related to E or H, instead of solving all PDEs related to both E and H. For 2-D MT forward modeling, the conductivity varies only in the vertical direction $z$ and the horizontal direction $y$; therefore, we can decouple equation 2 into two modes: mode $xy(E_{x},H_{y},H_{z})$ and mode $yx(H_{x},E_{y},E_{z})$ by
@@
where $u_{bc}$ presents the electric or magnetic field at the boundary, and $g(y,z)$ denotes the Dirichlet boundary condition at the coordinates $y$ and $z$. Once we obtain E and H, we can determine the apparent resistivity $\rho_{xy}$ and $\rho_{yx}$, as well as the phase $\phi_{xy}$ and $\phi_{yx}$, using
@@
where Re denotes the real part and Im denotes the imaginary part, respectively. Additionally, our primary concern is the electromagnetic field on the Earth surface, as this is where we are typically able to deploy instruments for the collection of electromagnetic field data. To obtain the apparent resistivity and phase on the surface, conventional numerical modeling methods, such as FDM, are often used to solve equation 3, which can be represented as
@@
where $G_{\theta}$ indicates the solver with parameter $\theta$ for the desired apparent resistivity and phase. The classical linear MT forward modeling method, such as FDM, usually requires discretizing the conductivity, followed by constructing linear systems for different frequencies. Hence, the computational cost of conventional methods is very high as their performance depends on the number of grids, and they require constructing a separate operator for each frequency. Therefore, it is necessary to develop a surrogate model of $G_{\theta}$ to improve the efficiency of MT forward modeling.
@@ 2.2.1 Neural Operators
Neural operators have been used to solve PDEs with a modern machine learning technique, which is the solution operator. Representative models are FNOs (Li et al., 2020) and DeepONets (Lu et al., 2021).
To construct a DNN for solving PDEs, Li et al. (2020) proposed the NO, which can be written as
@@
where $v_{j+1}$ denotes the output of the $j+1$-th layer of the DNN, $\alpha$ represents the activation function, $v_{j}$ indicates the output of the previous layer, and $W$ is the weight. Compared to the standard DNN, the right-hand side of equation 7 includes an additional integral term, where $\kappa$ signifies the integral kernel function that must be learned from the data. By incorporating this integral term, the NO can extract non-local features from the data. Due to the fact that integral computation is time consuming, Li et al. (2021) further proposed the FNO by the Fourier layer that transforms the integral operation of equation 7 into frequency domain. The Fourier layer can be expressed as
@@
where $\mathcal{F}$ denotes the Fourier transform and $\mathcal{F}^{-1}$ is the inverse Fourier transform. As shown in Fig. 1, FNO consists of a lifting layer $P$, a Fourier layer, and a projecting layer $Q$. The lifting layer $P$ maps the input $a$ to a high-dimensional channel space. The Fourier layer transforms the input $v$ into frequency and filters the Fourier coefficients, followed by the inverse Fourier transform. The $W$ represents linear transformations, such as those performed by MLPs. The projecting layer $Q$ projects the data to the desired dimension.
Although FNO can share the same parameters regardless of the discretization used for both input and output (Kovachki et al., 2024), it is not suitable for MT forward modeling. This is because FNO typically requires that input and output data have consistent spatial dimensions. This means that if the input data is a multi-dimensional array with a specific resolution, then the output data should also have the same dimensions and resolution. For MT forward modeling, the results of the forward modeling depend not only on the resistivity model but also on the observation frequency and the location of the electromagnetic field observation.
Unlike FNO, DeepONet is not restricted to a specific architecture and can incorporate various types of neural network architectures into its branch and trunk nets, providing greater flexibility. The formulation of DeepONet can be expressed as
@@
where $b_{k}$ represents $k$-th branch net, $t_{k}$ denotes $k$-th trunk net, and $p$ is the number of branches. In DeepONet, branch and trunk nets facilitate the approximation of infinite-dimensional mappings from inputs to outputs, offering a highly flexible and efficient approach. The branch network provides the necessary input features, while the trunk network maps these features to the output space, allowing DeepONet to handle complex and high-dimensional operator learning tasks (Tianping Chen and Hong Chen, 1995). This architecture allows DeepONet to generalize well in a range of scenarios, including different domain geometries, input parameters, and initial and boundary conditions (Kontolati et al., 2024).
@@ 2.2.2 KANs
In KANs, the activation functions are not fixed; instead, they are learned during the training process, allowing for a more flexible and efficient approximation of complex functions. Each weight parameter in a KAN is replaced by a univariate function. In the original implementation (Liu et al., 2024), $\psi(x)$ is defined as a weighted combination of a basis function $b(x)$ and B-splines,
@@
where $b(x)$ and the spline$(x)$ are defined as follows.
@@
where $w_{b}$,$w_{s}$ and $c_{i}$ are the weights optimized during training, and $B_{i}(x)$ are the basis functions B-spline defined over a grid.
Thus, the proposed method is able to improve interpretability, as the learnable functions can be visualized and understood more intuitively. The training process involves optimizing the weights of these splines to minimize the loss function, adjusting the shape of the spline to best fit the training data. KANs offer a promising alternative to traditional MLPs by leveraging the theoretical foundation of the Kolmogorov-Arnold representation theorem, leading to potentially more efficient and interpretable models for approximating complex functions.
@@ 2.2.3 EFKAN
Considering the characteristics of MT forward modeling, as well as the advantages of FNOs, DeepONets, and KANs, we propose to extend FNO with KAN to improve the accuracy of MT forward modeling; we name it EFKAN (Fig. 2). Specifically, we employ FNO as the branch net to map the resistivity $\rho(y,z)$ to $U(y,z)$; for the trunk net, we utilize the desired frequencies and coordinates of the electromagnetic field observation as input to KAN to obtain $D(y,z,f)$. Note that $D(y,z,f)$ has the same dimension as $U(y,z)$. We can therefore achieve the desired apparent resistivity and phase ($\rho_{xy},\phi_{xy},\rho_{yx},\phi_{yx}$) by multiplying the matrix of $u_{i}$ and $d_{i}$. For the 2-D case, the mathematical expression of MT forward modeling with EFKAN can be written as
@@
In the FNO of EFKAN, we utilize the fast Fourier transformer (FFT), 2-D convolution, and GELU to construct the Fourier layer. In the frequency domain, it is necessary to truncate the high modes to obtain $\mathcal{F}\left(v_{t}\right)\in\mathbb{C}^{k_{\max}\times d_{v}}$ because we convolve $v_{t}\in\mathbb{R}^{n\times d_{v}}$ with a function that has only $k_{\max}$ Fourier modes. We set $k_{\max}$ to 18. For the 2-D convolution, we define the kernel size to $1\times 1$, and keep the input channels equal to the output channels. Additionally, we place linear layers at the beginning and end of the FNO to determine the dimension of output, and we employ the Gaussian error linear unit (GELU) as the activation function. More details related to FNO can be found in Table 1. For the KAN of EFKAN, we use the original KAN implemented in (Liu et al., 2024) that adopts the B-spline function as the activation function. The trunk net-KAN has three layers, and the details about each layer are listed in Table 2.